# OccluTrack: Rethinking Awareness of Occlusion for Enhancing Multiple Pedestrian Tracking

### v1

@@ Abstract
Multiple pedestrian tracking faces the challenge of tracking pedestrians in the presence of occlusion. Existing methods suffer from inaccurate motion estimation, appearance feature extraction, and association due to occlusion, leading to inadequate Identification F1-Score (IDF1), excessive ID switches (IDSw), and insufficient association accuracy and recall (AssA and AssR). We found that the main reason is abnormal detections caused by partial occlusion. In this paper, we suggest that the key insight is explicit motion estimation, reliable appearance features, and fair association in occlusion scenes. Specifically, we propose an adaptive occlusion-aware multiple pedestrian tracker, OccluTrack. We first introduce an abnormal motion suppression mechanism into the Kalman Filter to adaptively detect and suppress outlier motions caused by partial occlusion. Second, we propose a pose-guided re-ID module to extract discriminative part features for partially occluded pedestrians. Last, we design a new occlusion-aware association method towards fair IoU and appearance embedding distance measurement for occluded pedestrians. Extensive evaluation results demonstrate that our OccluTrack outperforms state-of-the-art methods on MOT-Challenge datasets. Particularly, the improvements on IDF1, IDSw, AssA, and AssR demonstrate the effectiveness of our OccluTrack on tracking and association performance.
@@ I Introduction
Multiple pedestrian tracking is a challenging task aiming to form trajectories for detected pedestrians over time. This process involves detecting pedestrians and associating identical pedestrians from sequential frames. It is important in various real-world applications such as surveillance [1, 2], robotics [3, 4], and autonomous driving [5, 6].
@@
Existing state-of-the-art multiple pedestrian tracking methods [7, 8, 9, 10] associate detected pedestrians by combining cues like motions and appearance features to address the occlusion problem. Kalman Filter, one of the motion estimators, is commonly used to provide motion cues by formulating multiple pedestrian tracker as a linear estimation problem, which recursively predicts and updates trajectories from noisy observations (detections) over frames. Re-ID modules extract appearance features from multiple frames to identify the same pedestrian. However, these methods, even with re-ID, are insufficient to resolve occlusion problems because of inaccurate motion estimates, unreliable appearance features, and unfair association.
@@
Recently, occlusion has drawn much attention in multiple pedestrian tracking, and various approaches were proposed to address this issue, e.g., tracking by attention [11, 12, 13], graph neural networks [14], self and cross-attention mechanism [15, 16, 17, 18], and hierarchical feature extraction [19]. They address the occlusion problem by formulating motion patterns via advanced deep learning models in adjacent or multiple frames [15, 16, 18, 11] or understanding the context information in the scene [19, 14, 17]. However, these methods require significant computation resources compared with the Kalman Filter, and they ignore the effects caused by partial occlusion.
@@
Figure 1: The motivation of our proposed OccluTrack. (a) Left: SOTA trackers struggle with abnormal motions caused by partial occlusion, leading to inaccurate motion estimation. Middle: They also lack pose guidance during partial occlusion, resulting in unreliable appearance features. Right: Treating occluded and non-occluded pedestrians equally in distance measurement hinders optimal association. (b) In contrast, OccluTrack incorporates abnormal motion suppression (left), pose-guided re-ID (middle), and occlusion-aware distance measurement (right) to address these limitations.
@@
Through visualization of experimental results, we found partial occlusion is the missing key to resolving the occlusion problem. Partial occlusion creates abnormal bounding boxes covering some body parts only. These abnormal bounding boxes change suddenly in center points and aspect ratios, which causes “inaccurate motion estimates”, appearance features, and associations. As shown in Fig. 1(a), abnormal detections caused by partial occlusion mislead the motion estimator in predicting wrong trajectories. Especially the errors will be accumulated and amplified without accurate observations during full occlusion. As for re-ID modules, the inputs are from the cropped images according to the detections. Nevertheless, the cropped images will cover two persons (part of the obscured and part of the visible person) during partial occlusion, introducing “unreliable and noisy appearance features”. Moreover, the predictions from the motion estimator during occlusion are not as accurate as those in normal situations because of the error accumulation caused by partial occlusion. Hence, the association method should not treat occluded persons as strictly as visible persons, i.e., “unfair association”.
@@
To alleviate the three problems caused by partial occlusions, we propose the OccluTrack, an adaptive occlusion-aware multiple pedestrian tracker with three strategies, as shown in Fig. 1(b). First, we propose an abnormal motion suppression mechanism for stabilizing parameter updates in the Kalman Filter. In particular, the abnormal motion suppression mechanism leverages the history of tracked persons’ observations to detect and suppress abnormal motions. The motion can be more accurately predicted by abnormal motion suppression when partial occlusion occurs. Second, we introduce a pose-guided re-ID strategy for robust part feature extraction. The pose-guided re-ID strategy utilizes a pose estimator to guide the feature extractor to obtain more discriminative features. Lastly, we adopt an occlusion-aware distance measurement for occluded person association. The occlusion-aware distance measurement strategy involves a combination of IoU distance and appearance embedding distance under adaptive thresholds based on the level of occlusion, which is fair for occluded pedestrians.
@@
We apply our method to three baseline trackers (JDE [20], FairMOT [8], and BoT-SORT[10]), showing that the IDF1 and IDSw on the MOT17 validation set are improved consistently from 1.5% to 2.4% and 15% to 30%, respectively. We also evaluated our OccluTrack on the MOTChallenge datasets, and the results show that our method outperforms other approaches by a significant margin. Compared with the state-of-the-art BoT-SORT [10], the average improvement on IDF1 and IDSw reaches 1.0% and 12.5% on MOT17 and MOT20 test sets. Meanwhile, the improvements in AssA and AssR show a better association ability compared with other methods. Especially on the MOT20 dataset, AssA and AssR increase by 2.7% and 1.7% respectively. Our contributions can be summarized as follows:
@@
    We propose the Kalman Filter with an abnormal suppression mechanism to explicitly alleviate abnormal motions caused by partial occlusion.
@@
    Our proposed OccluTrack outperforms the state-of-the-art methods on MOTChallenge datasets, showing better tracking and association performance.
@@ II-A Multiple Pedestrian Tracking
At the early stage, works on multiple pedestrian tracking mainly focus on solving occlusion problems by motion cues. Intuitively, IoUTrack [21] associates the tracklets and detection solely based on the IoU. Only the IoU distance for detections from two adjacent frames indicates the geometric similarity. SORT [7] was the first work introducing Kalman Filter as the motion estimator into tracking-by-detection methods. Unlike IoUTrackers, SORT can predict the current positions upon the motion cues generated from Kalman Filter and enhance the tracking performance during occlusion. ByteTrack [9] and OC-SORT [22] promoted the SORT by associating almost every detected object and smoothing the Kalman Filter by observations. Apart from SORT-based methods, CenterTrack [23] built on CenterNet [24] adopts a regressor as the motion estimator, which regresses object motions from two adjacent frames. Following works [25, 26, 27, 28] make CenterTrack more robust by introducing more proper motion formulation methods like incorporating past detections, cost volume, and Quasi-Dense similarities. STTD [29] detects pedestrians with topologies and learns the dynamics of moving pedestrians within the same topology group.
@@
Motion cues are not enough for occlusion avoidance because of a lack of accurate observations. The appearance feature [30] is another cue to address the occlusion problem. DeepSort [31] first introduced re-ID to solve the occlusion using appearance features. Following works [20, 8, 32, 33, 10, 34, 35, 36] adopt embedded or stand-by more robust re-ID modules to enhance the ability to resolve the occlusion problem by modeling re-ID more accurately. However, appearance features vary in different scenarios, especially during partial occlusion.
@@
Besides, some tracking-by-attention [11, 13, 12, 37, 38] methods modify DETR [39, 40] to an end-to-end tracker by introducing tracking queries and detection queries for tracked objects and new-born objects. These tracking-by-attention methods lack motion cues, which are sensitive to occlusions. Methods like P3AFormer [41] re-introduce motion cues by methods like optical flows. Some other methods solve the occlusion problem from aspects of learning long-term temporal features and understanding scenes. DeepTracklet and MotionTrack [15, 16, 17, 18] make use of self- or cross-attention to learn long-term features. TransMOT [14] and Quo Vadis[19] try to understand scenes from the perspective of graphs and Bird-Eye-View. However, these methods usually take dramatic computation resources and ignore the effects caused by partial occlusion.
II-B Kalman Filter
@@
Kalman Filter is commonly used in multiple pedestrian tracking with a data association algorithm such as the nearest neighbor and Hungarian algorithms. Data association algorithms associate the observations with the predicted states, and the Kalman Filter updates the state estimates based on the associations. It is a recursive algorithm that estimates the state of a system based on noisy observations. It consists of two main steps: the prediction step and the update step. The two predictions are given by Eq.(1). At each time step $t$, the Kalman Filter first predicts the state of the system $\hat{x}_{t|t-1}$ using the previous state estimate $\hat{x}_{t-1|t-1}$. And matrix $A$ is the state-transition matrix that relates the state at time $t-1$ to the state at time $t$. Additionally, the Kalman Filter also predicts the covariance of the state estimate $P_{t|t-1}$ using the previous covariance estimate $P_{t-1|t-1}$ and a process noise $Q$ that describes how the uncertainty in the state evolves.
$\text{Predict}:\begin{cases}\hat{x}_{t|t-1}=A\hat{x}_{t-1|t-1}\\ P_{t|t-1}=AP_{t-1|t-1}A^{T}+Q\end{cases}.$ (1)
@@
After predicting the state and covariance estimates, the Kalman Filter updates them based on the current observation (or measurement) $z_{t}$, as formulated by Eq.(2). The observation model relates the observation to the state $z_{t}=Hx_{t}+v_{t}$, where $H$ is the observation matrix that maps the state to the observation space, and $v_{t}$ is the observation noise with the covariance matrix of $R$. The Kalman gain $K_{t}$ is first calculated in the update step, which determines the weight given to the observation. The updated state estimate $x_{t|t}$ and covariance estimate $P_{t|t-1}$ are then computed. $I$ is the identity matrix.
$\text{Update}:\begin{cases}K_{t}=P_{t|t-1}H^{T}(HP_{t|t-1}H^{T}+R)^{-1}\\ \hat{x}_{t|t}=\hat{x}_{t|t-1}+K_{t}(z_{t}-H\hat{x}_{t|t-1})\\ P_{t|t}=(I-K_{t}H)P_{t|t-1}\end{cases}.$ (2)
@@ III-A Overview
Figure 2: Overview of our proposed tracker. First, we utilize YOLOX [42] as the object detector to obtain bounding boxes and separate them into high-score and low-score detections. For high-score detections, the re-ID module is used to extract appearance features. Second, the association module is to associate the tracklets of previous frames with the high-score detections (with feature embeddings) and low-score detections of current frames sequentially. Finally, the Kalman Filter (KF) is updated based on the matching results obtained from the association step. Additionally, the appearance features are updated based on the matching results. We highlight three main contributions (components) for MOT: abnormal motion suppression of KF in (c), pose-guided re-ID in (a), and occlusion-aware distance measurement in (b).
Fig. 2 illustrates the overview of our proposed OccluTrack built upon the baseline tracker, BoT-SORT [10]. To enhance the tracking performance during occlusion, we introduce three components: abnormal motion suppression-based Kalman Filter, pose-guided re-ID, and occlusion-aware distance measurement. For the update component in Fig. 2(c), we propose an abnormal motion suppression mechanism in the Kalman Filter. It first detects abnormal motions by the difference between the current speed and a residual speed from a speed filter. Based on the difference, it determines a dynamic suppression parameter to suppress the outlier motions in the update step of the Kalman Filter. For the re-ID component in Fig.2(a), our OccluTrack incorporates a pose-guided re-ID module for robust appearance feature extraction. Our pose-guided re-ID extracts local and global features with pose guidance and combines them with an adaptive fusion operation. For the association component in Fig. 2(b), we introduce an occlusion-aware distance measurement to facilitate fair and accurate association of occluded persons by leveraging an adaptive threshold based on occlusion situations. We will elaborate on them in the following sections.
We present the tracking algorithm for our proposed OccluTrack, as shown in Algorithm 1. The algorithm is based on BoT-SORT [10] and adopts the proposed pose-guided re-ID (PG_ReID), occlusion-aware distance measurement (ODM), and Kalman Filter with abnormal motion suppression (KF_AMS) modules indicated by the green comments. The algorithm begins with initializing the tracklets $T=[T_{b},T_{E},BF_{B}]$ with [,,][$\emptyset$,$\emptyset$,$\emptyset$][ ∅ , ∅ , ∅ ] for bounding boxes, feature embeddings, and buffer of the tracked bounding boxes. Next, each frame from a video clip is processed sequentially. For every frame $v$ in video clip $V$, we perform a pedestrian detection method (Det) to obtain the bounding boxes $B$. These bounding boxes are divided into high-score detections $B_{h}$ and low-score detections $B_{l}$ based on their confidence scores. For $B_{h}$, we extract the corresponding regions from the frame and feed them into the PG_ReID module, as shown in Fig.2(a), to compute the feature embeddings $E$. In the prediction of Kalman Filter, we predict the bounding boxes $T_{B}$ for the tracklets in the current frame and perform motion compensation as BoT-SORT [10]. In data association, illustrated in Fig.2(b), we calculate the distance matrix $\hat{d}_{h}$ via the ODM for high-score detections with embeddings. The Hungarian algorithm is then employed to find the indices for matched tracklets and detections, as well as unmatched tracklets and detections. For low-score detections and unmatched tracklets from the previous step, we compute the distance matrix $\hat{d}_{l}$ using the IoU distance metric to determine the matched and unmatched indices by the Hungarian algorithm. As shown in Fig.2(c), we update KF_AMS parameters and bounding box buffer $BF_{B}$ for the tracklets with matched high-score detections and embeddings, as well as update KF_AMS parameters and bounding box buffer $BF_{B}$ for the tracklets with matched low-score detections according to matched indices. Remained tracklets that are untracked for a duration exceeding a threshold are removed, while new tracklets are initialized for remaining detections $B^{remain}$ if their confidence score exceeds a threshold $\theta_{new\_track}$.

### v2

@@ Abstract
Multiple pedestrian tracking is crucial for enhancing safety and efficiency in intelligent transport and autonomous driving systems by predicting movements and enabling adaptive decision-making in dynamic environments. It optimizes traffic flow, facilitates human interaction, and ensures compliance with regulations. However, it faces the challenge of tracking pedestrians in the presence of occlusion. Existing methods overlook effects caused by abnormal detections during partial occlusion. Subsequently, these abnormal detections can lead to inaccurate motion estimation, unreliable appearance features, and unfair association. To address these issues, we propose an adaptive occlusion-aware multiple pedestrian tracker, OccluTrack, to mitigate the effects caused by partial occlusion. Specifically, we first introduce a plug-and-play abnormal motion suppression mechanism into the Kalman Filter to adaptively detect and suppress outlier motions caused by partial occlusion. Second, we develop a pose-guided re-identification (Re-ID) module to extract discriminative part features for partially occluded pedestrians. Last, we develop a new occlusion-aware association method towards fair Intersection over Union (IoU) and appearance embedding distance measurement for occluded pedestrians. Extensive evaluation results demonstrate that our method outperforms state-of-the-art methods on MOTChallenge and DanceTrack datasets. Particularly, the performance improvements on IDF1 and ID Switches, as well as visualized results, demonstrate the effectiveness of our method in multiple pedestrian tracking.
@@ I Introduction
Multiple pedestrian tracking is indispensable for intelligent transport [1] and autonomous driving systems [2, 3] as it enables predicting pedestrian movements [4], adaptive decision-making [5, 6], and compliance with regulations [7] in dynamic urban environments. By accurately monitoring pedestrian behavior, this technology enhances safety by minimizing collision risks [8] and optimizing traffic flow [9]. Additionally, it facilitates efficient navigation for autonomous vehicles through crowded areas and ensures smooth interactions between vehicles and pedestrians. This task remains challenging, aiming to form trajectories for detected pedestrians over time. This process involves detecting pedestrians and associating identical pedestrians from sequential frames.
@@
Existing state-of-the-art multiple pedestrian tracking methods [10, 11, 12, 13] associate detected pedestrians by combining cues like motions and appearance features to address the occlusion problem. Kalman Filter, one of the motion estimators, is commonly used to provide motion cues by formulating multiple pedestrian trackers as a linear estimation problem, which recursively predicts and updates trajectories from noisy observations (detections) over frames. Re-ID modules extract appearance features from multiple frames to identify the same pedestrian. However, these methods, even with Re-ID, are insufficient to resolve occlusion problems because of inaccurate motion estimates, unreliable appearance features, and unfair association.
@@
Recently, occlusion has drawn much attention in multiple pedestrian tracking, and various approaches were proposed to address this issue, e.g., tracking by attention [14, 15], graph neural networks [16], self and cross-attention mechanism [17, 18], and hierarchical feature extraction [19]. They address the occlusion problem by formulating motion patterns via advanced deep learning models in adjacent or multiple frames or understanding the context information in the scene. However, these methods require significant computation resources compared with the Kalman Filter, and they ignore the effects caused by partial occlusion.
@@
Figure 1: The motivation of our proposed OccluTrack. (a) Left: the SOTA tracker [13] struggles with abnormal motions caused by partial occlusion, leading to inaccurate motion estimation. Middle: They also lack pose guidance during partial occlusion, resulting in unreliable appearance features. Right: Treating occluded and non-occluded pedestrians equally in distance measurement hinders optimal association. (b) In contrast, OccluTrack incorporates abnormal motion suppression (left), pose-guided Re-ID (middle), and occlusion-aware distance measurement (right) to address these limitations.
@@
Through the visualization of experimental results, we found that partial occlusion is the missing key to resolving the occlusion problem. Partial occlusion creates abnormal bounding boxes covering some body parts only. These abnormal bounding boxes change suddenly in center points and aspect ratios, which causes inaccurate motion estimates, appearance features, and associations. As shown in Fig. 1(a), abnormal detections caused by partial occlusion mislead the motion estimator in predicting wrong trajectories. Especially the errors will be accumulated and amplified without accurate observations during full occlusion. As for Re-ID modules, the inputs are from the cropped images according to the detections. Nevertheless, the cropped images will cover two persons (part of the obscured and part of the visible person) during partial occlusion, introducing “unreliable and noisy appearance features”. Moreover, the predictions from the motion estimator during occlusion are not as accurate as those in normal situations because of the error accumulation caused by partial occlusion. Hence, the association method should not treat occluded persons as strictly as visible persons, i.e., “unfair association”.
@@
To alleviate the effects caused by partial occlusions in motion estimation, Re-ID, and association, we propose the OccluTrack, an adaptive occlusion-aware multiple pedestrian tracker with three strategies, as shown in Fig. 1(b). First, we propose a plug-and-play abnormal motion suppression mechanism for stabilizing parameter updates in the Kalman Filter without effects on inference speed. In particular, the abnormal motion suppression mechanism leverages the history of tracked observations to detect and suppress abnormal motions. The motion can be more accurately predicted by abnormal motion suppression when partial occlusion occurs. Second, we introduce a pose-guided Re-ID strategy for robust part feature extraction. The pose-guided Re-ID strategy utilizes a pose estimator to guide the feature extractor to obtain more discriminative features. Lastly, we adopt an occlusion-aware distance measurement for occluded person association. The occlusion-aware distance measurement strategy involves a combination of IoU distance and appearance embedding distance under adaptive thresholds based on the level of occlusion, which is fair for occluded pedestrians.
@@
We apply our method to three baseline trackers (JDE [20], FairMOT [11], and BoT-SORT[13]), showing that the IDF1 and IDSw on the MOT17 validation set are improved consistently from 1.5% to 2.4% and 15% to 30%, respectively. We also evaluated our OccluTrack on the MOTChallenge [21, 22] and DanceTrack [23] datasets, and the results show that our method outperforms other approaches by a significant margin under both private and public protocols.
@@
    We propose an abnormal suppression mechanism in the Kalman Filter to explicitly alleviate abnormal motions caused by partial occlusion.
@@
    Extensive experiments were conducted on MOTChallenge and DanceTrack datasets, showing our proposed OccluTrack outperforms the state-of-the-art methods and has better tracking and association performance.
@@ II-A Multiple Pedestrian Tracking
At the early stage, works on multiple pedestrian tracking mainly focus on solving occlusion problems by motion cues. Intuitively, IoUTrack [24] associates the tracklets and detection solely based on the IoU. Only the IoU distance for detections from two adjacent frames indicates the geometric similarity. SORT [10] was the first work introducing Kalman Filter as the motion estimator into tracking-by-detection methods. Unlike IoUTrackers, SORT can predict the current positions upon the motion cues generated from Kalman Filter and enhance the tracking performance during occlusion. ByteTrack [12] and OC-SORT [25] promoted the SORT by associating almost every detected object and smoothing the Kalman Filter by observations. Apart from SORT-based methods, CenterTrack [26] built on CenterNet [27] adopts a regressor as the motion estimator, which regresses object motions from two adjacent frames. The following works [28, 29] make CenterTrack more robust by introducing more proper motion formulation methods like incorporating past detections, cost volume, and Quasi-Dense similarities. STTD [30] detects pedestrians with topologies and learns the dynamics of moving pedestrians within the same topology group.
@@
Motion cues are not enough for occlusion avoidance because of a lack of accurate observations. The appearance feature [31] is another cue to address the occlusion problem. DeepSort [32] first introduced Re-ID to solve the occlusion using appearance features. Following works [11, 13, 33] adopt embedded or stand-by more robust Re-ID modules to enhance the ability to resolve the occlusion problem by modeling Re-ID more accurately. However, appearance features vary in different scenarios, especially during partial occlusion.
@@
Besides, some tracking-by-attention [14, 34] methods modify DETR [35, 36] to an end-to-end tracker by introducing tracking queries and detection queries for tracked objects and new-born objects. These tracking-by-attention methods lack motion cues, which are sensitive to occlusions. Methods like P3AFormer [37] re-introduce motion cues by methods like optical flows. Some other methods solve the occlusion problem from aspects of learning long-term temporal features and understanding scenes. DeepTracklet and MotionTrack [17, 18] make use of self- or cross-attention to learn long-term features. TransMOT [16] and Quo Vadis[19] try to understand scenes from the perspective of graphs and Bird-Eye-View. However, these methods usually take dramatic computation resources and ignore the effects caused by partial occlusion.
II-B Motion Estimation with Kalman Filter
@@
Kalman Filter is commonly used in multiple pedestrian tracking with a data association algorithm such as the nearest neighbor and Hungarian algorithms. Data association algorithms associate the observations (e.g., object detections) with the predicted states, and the state estimates of the Kalman Filter are updated based on the associations. It is a recursive algorithm that estimates the state of a system based on noisy observations. It consists of two main steps: the prediction step and the update step. The two predictions are given by Eq. (1). At each time step $k$, the Kalman Filter first predicts the state of the system $\hat{s}_{k|k-1}$ using the previous state estimate $\hat{s}_{k-1|k-1}$. And matrix $A$ is the state-transition matrix that relates the state at time $k-1$ to the state at time $k$. Additionally, the Kalman Filter also predicts the covariance of the state estimate $P_{k|k-1}$ using the previous covariance estimate $P_{k-1|k-1}$ and a process noise $Q$ that describes how the uncertainty in the state evolves.
$\text{Predict}:\begin{cases}\hat{s}_{k|k-1}=A\hat{s}_{k-1|k-1}\\ P_{k|k-1}=AP_{k-1|k-1}A^{\top}+Q\end{cases}.$ (1)
@@
After predicting the state and covariance estimates, the Kalman Filter updates them based on the current observation (or measurement) $z_{k}$, as formulated by Eq. (2). The observation model relates the observation to the state $z_{k}=Hs_{k}+v_{k}$, where $H$ is the observation matrix that maps the state to the observation space, and $v_{k}$ is the observation noise with the covariance matrix of $R$. The Kalman gain $G_{k}$ is first calculated in the update step, which determines the weight given to the observation. The updated state estimate $\hat{s}_{k|k}$ and covariance estimate $P_{k|k-1}$ are then computed. $I$ is the identity matrix.
$\text{Update}:\begin{cases}G_{k}=P_{k|k-1}H^{\top}(HP_{k|k-1}H^{\top}+R)^{-1}\\ \hat{s}_{k|k}=\hat{s}_{k|k-1}+G_{k}(z_{k}-H\hat{s}_{k|k-1})\\ P_{k|k}=(I-G_{k}H)P_{k|k-1}\end{cases}.$ (2)
II-C Appearance Feature Extraction
Appearance-based methods have proven to be effective in addressing occlusion challenges within the realm of multiple pedestrian tracking. Person re-identification (Re-ID) is a thoroughly researched problem in computer vision, with numerous studies proposing innovative algorithms and techniques to tackle it. In existing research [38, 39, 40], appearance features are typically modeled using both global and local features. Global features encompass information extracted from the entire body, whereas local features are derived from various body parts. However, these conventional approaches may not adequately adapt to the demands of multiple pedestrian tracking, particularly in terms of efficiency and accounting for changes in appearance over time. In the domain of multiple pedestrian tracking, most existing methods primarily rely on global features, which can be susceptible to issues related to occlusion. For instance, DeepSORT [32] and subsequent methods such as BoT-SORT [13], StrongSORT [33], and Deep OC-SORT [41] utilize a pre-existing Re-ID module to extract global features. Similarly, FairMOT [11] and its subsequent trackers [42, 43] employ feature embeddings obtained from the detection feature map to represent global appearance. While some recent methods [44, 45] recognize the importance of incorporating local features to address occlusion challenges, they often employ complex network architectures like transformers, which may not be conducive to real-time tracking efficiency.
II-D Association for Occlusion
The Hungarian algorithm plays a crucial role in linking detections and tracklets in tracking-by-detection methods by incorporating both motion and appearance distances. In the early stages of tracking-by-detection methods [10, 26], the motion distance for association was typically computed using metrics such as Intersection over Union (IoU) or center distance. However, as the importance of appearance features became evident, the concept of appearance distance was introduced alongside motion distance to improve tracking performance, especially in scenarios involving occlusions. To optimize association in the presence of occlusions, various distance measurement strategies based on motion and appearance distances have been proposed. Most methods employ a weighted distance measurement approach [46], where the two distances are combined using weighted averaging. More recently, approaches like BoT-SORT [13] and Deep OC-SORT [41] have introduced innovative measurement methods. These methods first mask out the appearance distance using the motion distance and then select the minimum distance between them for the association. However, both weighted distance measurement and masked distance measurement tend to overlook the effects caused by partial occlusion. Partial occlusion can lead to unreliable predictions during occlusion periods. Consequently, applying the same distance measurement requirements for both occluded and visible pedestrians may not be equitable or optimal in addressing the challenges posed by partial occlusion.
@@ III-A Overview
Figure 2: Overview of our proposed tracker. First, we utilize YOLOX [47] as the object detector to obtain bounding boxes and separate them into high-score and low-score detections. For high-score detections, the pose-guided Re-ID module in (a) is used to extract appearance features. Second, the abnormal motion suppression Kalman Filter (AMS KF) is proposed to obtain $Predictions_{k-1}$ from previous $tracklets_{k-1}$. Based on the predictions, the association module with occlusion-aware distance measurement in (b) first associates the $Predictions_{k-1}$ of previous frames with the high-score detections (with feature embeddings), and then the second association matches unmatched predictions with low-score detections. Finally, the abnormal motion suppression Kalman Filter in (c) is updated based on the matching results obtained from the association step. Specifically, the proposed AMS KF first detects the abnormal detections with abnormal changes. The $tracklets_{k}$ and parameters of the KF are updated correspondingly by considering the outlier detections caused by partial occlusion.
Fig. 2 illustrates the overview of our proposed OccluTrack built upon the baseline tracker, BoT-SORT [13]. It efficiently handles video frame processing for improved tracking in occlusion scenarios. In OccluTrack, it first performs object detections and separates detections into high-score and low-score detections. For high-score detections, we utilize our proposed pose-guided Re-ID module in Fig. 2(a) to extract appearance features. Afterward, the abnormal motion suppression Kalman Filter (AMS KF) is used to provide predictions based on the tracklets from previous frames and pass the predictions to the occlusion-aware association module. After that, detected pedestrian bounding boxes and appearance features are used to associate with detection and predictions from the Kalman Filter. Especially, the high-score detections are associated with our developed occlusion-aware distance measurement in Fig. 2(b). For low-score detections, we associate them with predictions that are not matched with high-score detections using IoU distance. And associated or matched detections are used to update our abnormal motion suppression Kalman Filter in Fig. 2(c). During the update, our proposed AMS KF will detect abnormal motions and suppress them when updating AMS KF parameters.
We present the detailed tracking algorithm for our proposed OccluTrack, as shown in Algorithm 1. The tracking process is enhanced with specialized modules: pose-guided Re-ID (PG_ReID), occlusion-aware distance measurement (ODM), and an abnormal motion suppression Kalman Filter (AMS_KF). The algorithm starts by initializing tracklets with empty sets for bounding boxes, feature embeddings, and tracked bounding boxes buffer. For each frame in a video, pedestrian detection identifies bounding boxes, which are then categorized into high and low-confidence detections. High-confidence detections undergo feature embedding extraction via the PG_ReID module. Kalman Filter predictions are made for tracklets, incorporating motion compensation. The distance matrix for high-confidence detections is calculated using ODM, and the Hungarian algorithm matches tracklets and detections. For low-confidence detections, matching is determined using the IoU distance metric, matched tracklets are updated with AMS_KF parameters and the bounding box buffer. Tracklets that remain untracked beyond a threshold are discarded, and new tracklets are initiated for remaining detections with confidence above a certain threshold. This comprehensive algorithm optimizes tracking performance, particularly in challenging occlusion scenarios.
III-B Abnormal Motion Suppression Kalman Filter
Kalman Filter has been widely adopted in tracking-by-detection methods, showing promising tracking performance. However, occlusion handling is a challenge within Kalman Filter. We observe that partial occlusion is one of the factors causing abnormal motions and wrong predictions during full occlusion, which is ignored in existing work. To mitigate the effects caused by partial occlusion, we introduce an abnormal motion suppression Kalman Filter to detect and suppress abnormal motions caused by partial occlusions during updating KF parameters.
Partial Occlusion in Kalman Filter
In the update step of the Kalman Filter, we find that the observation $z_{k}$ at time step $k$ is incorrect during partial occlusion due to abnormal detections, i.e., objects’ bounding boxes suddenly become small due to partial occlusion. We assume that $z_{k}$ is corrupted by an error $e_{k}$, i.e., $z^{\prime}_{k}=z_{k}+e_{k}$, where $z^{\prime}_{k}$ is the corrupted observation. Hence, the updated state estimate in Eq.(2) becomes:
$\tilde{s}_{k|k}=\hat{s}_{k|k-1}+G_{k}(z_{k}+e_{k}-H\hat{s}_{k|k-1})=\hat{s}_{k|k}+G_{k}e_{k},$ (3)
where $\tilde{s}_{k|k}$ is the corrupted state estimate disturbed by the $G_{k}e_{k}$, and $s=[x,y,w,h,\dot{x},\dot{y},\dot{w},\dot{h}]^{\top}$ denotes the state vector, which includes the bounding box position $[x,y,w,h]^{\top}$ and its moving velocity $[\dot{x},\dot{y},\dot{w},\dot{h}]^{\top}$. During partial occlusion, the error $e^{\prime}_{k}=G_{k}e_{k}=[e^{\prime}_{xk},e^{\prime}_{yk},e^{\prime}_{wk},e^{\prime}_{hk},e^{\prime}_{\dot{x}k},e^{\prime}_{\dot{y}k},e^{\prime}_{\dot{w}k},e^{\prime}_{\dot{h}k}]^{\top}$ will dominate the update process. Subsequently, this error will accumulate during full occlusion without observations. In this period, the Kalman Filter’s state update only relies on the predictions from the Kalman Filter. Here, we denote the duration for the full occlusion as $\tau$. As Eq. (1) indicates, the state is predicted by $\hat{s}_{k|k-1}=A\hat{s}_{k-1|k-1}$, where $A$ is the $n\times n$ transition matrix for the state from $k-1$ to $k$. Formally, $A$ can be elaborated for the linear motion as
When the person reappears after occlusion of $\tau$ from $k0$, with only the prediction of the Kalman Filter, the estimate can be predicted like :
$\begin{split}\tilde{s}_{k0+\tau}&=A^{\tau}\tilde{s}_{k0}=A^{\tau}(\hat{s}_{k0}+G_{k0}e_{k0})\\ &=\hat{s}_{k0+\tau}+A^{\tau}G_{k0}e_{k0}=\hat{s}_{k0+\tau}+A^{\tau}e^{\prime}_{k0}.\end{split}$ (5)
According to the transition matrix $A$ and error $e^{\prime}_{k}$, we can obtain $A^{\tau}$ and $e^{\prime}_{k0}$. Hence, the accumulated error $A^{\tau}e^{\prime}_{k0}$ can be calculated by
$A^{\tau}e^{\prime}_{k0}=\begin{bmatrix}1&0&0&0&\tau&0&0&0\\ 0&1&0&0&0&\tau&0&0\\ 0&0&1&0&0&0&\tau&0\\ 0&0&0&1&0&0&0&\tau\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&1\\ \end{bmatrix}\begin{bmatrix}e^{\prime}_{x0}\\ e^{\prime}_{y0}\\ e^{\prime}_{w0}\\ e^{\prime}_{h0}\\ e^{\prime}_{\dot{x}0}\\ e^{\prime}_{\dot{y}0}\\ e^{\prime}_{\dot{w}0}\\ e^{\prime}_{\dot{h}0}\\ \end{bmatrix}=\begin{bmatrix}e^{\prime}_{x0}+\tau e^{\prime}_{\dot{x}0}\\ e^{\prime}_{y0}+\tau e^{\prime}_{\dot{y}0}\\ e^{\prime}_{w0}+\tau e^{\prime}_{\dot{w}0}\\ e^{\prime}_{h0}+\tau e^{\prime}_{\dot{h}0}\\ e^{\prime}_{\dot{x}0}\\ e^{\prime}_{\dot{y}0}\\ e^{\prime}_{\dot{w}0}\\ e^{\prime}_{\dot{h}0}\\ \end{bmatrix}$ (6)
From Eq.(6), we can observe that the bounding box $[x,y,w,h]^{\top}$ is shifted by a fixed error $[e^{\prime}_{x0},e^{\prime}_{y0},e^{\prime}_{w0},e^{\prime}_{h0}]^{\top}$ and an accumulated error $[\tau e^{\prime}_{\dot{x}0},\tau e^{\prime}_{\dot{y}0},\tau e^{\prime}_{\dot{w}0},\tau e^{\prime}_{\dot{h}0}]^{\top}$ with a factor of $\tau$.